OpenStudy (anonymous):

a^6-b^3 I got a^6-b^3 Am I right????

OpenStudy (mathmath333):

what do u think

jhonyy9 (jhonyy9):

i think you need factoriz it ,yes ? so than you can using the formula a^3 -b^3 = ? hope you have learned about this

jhonyy9 (jhonyy9):

in this way : a^6 -b^3 = (a^2)^3 -b^3 = ... hope this will help you

OpenStudy (anonymous):

YA.